A father walks 36 miles in 12 hours, while his son covers the same distance on a bicycle at twice the father's rate of speed. At this rate how many miles would the bicycle rider travel in 9 hours?
Answer by checkley77(12844)

36/12=3 mph. is the father's walking rate.
2*3=6 mph. is the son's biking rate.
6*9=54 miles is the distance the son can bike in 9 hours.